A homeowner finds that there is a 0.15 probability that a flashlight does not work when turned on. If she has three flashlights, find the probability that at least one of them works when there is a power failure. Find the probability that the second flashlight works given that the first flashlight works.
Answer #1

1)P( at least one of them works when there is a power failure ) =1-P(all three does not work)

=1-(0.15)3 =0.996625

2)

since event of working of flashlights are independent:

probability that the second flashlight works given that the first flashlight works =P(flashlight works)

=1-0.15 =0.85